OpenStudy (anonymous):

If f(n) = n^2 - n then f(-4) is -20 20 12 -12

myininaya (myininaya):

Let's try to show the asker how to solve/evaluate problems like this, not just give the answer please.

OpenStudy (anonymous):

f(n) = n^2 - n (4) = (-4)^2 + 4 = 20

OpenStudy (anonymous):

The mods on here make it so uncomfortable to be free to do something. =\